### v2

# Transformer Block Coupling and its Correlation with Generalization in LLMs

Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices. By examining the relationships between these Jacobians, we uncover a transformer block coupling phenomenon in a variety of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension. We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories. These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
1 Introduction
In recent years, many openly available Large Language Models (LLMs) have been released, achieving state-of-the-art results on task-specific benchmarks. The abundance of models, each with differing architecture and training methodology, motivates comparing the underlying mechanisms that drive generalization.
Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023). The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; coupled refers to the interdependent token trajectories enabled by self-attention, and dynamical is due to the residual connections spanning various layers.
Throughout this work, we investigate whether we can identify structural characteristics that consistently occur in pretrained LLMs that are related to their generalization, as measured on the HuggingFace OpenLLM Leaderboard benchmark. To this end, we study and quantify the trajectories of token embeddings and their interactions, through linearizations of LLM layers given by the Jacobians of the associated transformer blocks.
This perspective raises several questions:
Q1. What regularity properties do these trajectories exhibit individually, and what are their relations with one another? More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
Q2. How do the properties of hidden representations and their relations emerge with training?
Q3. Are any of these properties related to the generalization capabilities of LLMs?
Figure 1: Relationship between Coupling and Benchmark Score. The plot illustrates the relationship between average coupling (negative mis-coupling) and benchmark score across various base LLMs (that are not fine-tuned). The data shows a noticeable trend: increased alignment correlates with improved benchmark scores. The regression fit produces an $R^{2}$ value of $0.7214$.
1.1 Contributions
We investigate the motivating questions across several openly-available LLMs, most having over 1B parameters, trained by 6 independent organizations, with varying training methods and data (Appendix A.1). Through our experiments, we identify consistent patterns that define the transformer block coupling phenomenon.
1. Coupling. The singular vectors of the Jacobians of transformer blocks couple across depth (Figures LABEL:fig:paris_all_8to16, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix). Further, coupling across Jacobians emerges with training (Figures 2, LABEL:fig:paris_all_8to16, LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
2. Generalization. The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1, LABEL:fig:alignment_vary_k). Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
3. Regularity. Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, 13, 15, 23, 24). Exponential growth occurs in contiguous token representations as a function of depth (Figures 14, 22, 23, LABEL:figs:expo-training).
We provide a new perspective on token embedding interactions within LLMs by examining transformers layers through their Jacobians. Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
Figure 2: Emergence of Singular Vector Coupling with Training. Normalized coupling (Section 3.1) evaluated for $J_{i,j}^{d_{1}},J_{i,j}^{d_{2}}$, with $|d_{1}-d_{2}|\leq 5$. Measurements are taken for checkpoints of Pythia 12B, Pythia 6.9B, Pythia 1B (Biderman et al., 2023) at training steps $\{1,16,128,13\text{k},23\text{k},33\text{k},\ldots,143\text{k}\}$. Coupling is low at early training iterations, and increases with training, with Pythia 12B and 6.9B displaying greater coupling than Pythia 1B in final training steps.
2 Background on Large Language Models
We describe LLMs as a deep composition of functions that iteratively transform token embeddings. In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked, they form a matrix:
$X^{0}=(x_{1}^{0},x_{2}^{0},\ldots,x_{n}^{0})\in\mathbb{R}^{n\times d_{\text{model}}}.$ (1)
The embeddings then pass through $L$ transformer blocks:
$X^{0}\xrightarrow{F_{\text{block}}^{1}}X^{1}\xrightarrow{F_{\text{block}}^{2}}\cdots X^{L-1}\xrightarrow{F_{\text{block}}^{L}}X^{L}.$ (2)
$X^{l}=F_{\text{block}}^{l}(X^{l-1})$ denotes the embeddings after the $l^{\text{th}}$ block, consisting of causal multi-headed attention (MHA), a feed-forward network (FFN), and normalization layers (LN) with residual connections:
$h^{l+1}(X^{l})$ $=\text{MHA}(\text{LN}(X^{l}))$ (3)
$g^{l+1}(X^{l})$ $=\text{LN}(X^{l}+h^{l+1}(X^{l}))$ (4)
$f^{l+1}(X^{l})$ $=h^{l+1}(X^{l})+\text{FFN}(g^{l+1}(X^{l}))$ (5)
$F_{\text{block}}^{l+1}(X^{l})$ $=X^{l}+f^{l+1}(X^{l}),$ (6)
where the MHA, LN, FFN are implicitly indexed by layer. Among many models (Appendix 1), an additional rotary positional embedding (RoPE, Su et al. (2023)) is applied in the MHA layer. In the final representation, typically an additional layer normalization is applied:
$F_{\text{block}}^{L}(X^{L-1})$ $=\text{LN}(X^{L-1}+h^{L}(X^{L-1})+\text{FFN}(g^{L}(X^{L-1}))).$ (7)
The output $X^{L}$ from the final block $F^{L}$ is passed into a bias-free linear layer $M\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{model}}}$, with $d_{\text{vocab}}$ denoting the size of the token vocabulary and $d_{\text{model}}$ is the dimension of the token embeddings. This layer $M$ computes final-layer logits for each token embedding, $\ell_{i}=Mx_{i}^{L}$. The prediction for the next token is then determined by selecting the maximal logit value: $\arg\max_{v\in\text{tokens}}\ell_{v,n}.$
3 Methods
3.1 Coupling of Singular Vectors of Jacobians
Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks given by their Jacobians
$J_{t_{1}t_{2}}^{l}=\frac{\partial}{\partial x_{t_{1}}^{l-1}}\left(f^{l}(X^{l-1})\right)_{t_{2}},$ (8)
defined for each layer $l\in\{1,\ldots,L\}$, and pair of tokens $t_{1},t_{2}\in\{1,\ldots,n\}$. Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023), but extended for cross-token interactions. Also, due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ if $t_{1}>t_{2}$, so we restrict our attention to the case $t_{1}\leq t_{2}$.
Singular value decomposition. The SVD of the Jacobians $J_{t_{1}t_{2}}^{l}$ are computed, i.e., $J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscripts $t_{1},t_{2}$, indicating the tokens, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
To measure coupling between Jacobians $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, define $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}$ for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, respectively.
Figure 3: Transformer Block Coupling. Displayed is a visualization of the various types of transformer block coupling, along with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$.
Depth-wise Coupling. To analyze coupling across transformer blocks, we fix $t$ (usually as the last token), and check alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through $A_{t}^{ll^{\prime},K}$ (we write the single subscript $t$ for clarity) across layers $l,l^{\prime}\in\{1,\ldots,L\}$.
Token-wise Coupling. To study the coupling across tokens, we start by fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, and analyze coupling across different cases of the $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}$. We start with the case where the input and output tokens are the same and compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which is the coupling across tokens for a token’s effect on its own trajectory. We also consider the context tokens’ impact on a trajectory by including coupling across cases with $t_{1}<t_{2}$. More specifically, we measure coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
Coupling measurement. To quantify the coupling of singular vectors of the Jacobians, we measure how far the $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ are from being diagonal. More precisely, we quantify the mis-coupling of $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ using $m:\mathbb{R}^{K\times K}\to\mathbb{R}$ defined by
$m(A)=\lVert A-\text{Diag}(A)\rVert_{F},$ (9)
where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. For normalized comparison between models across training in Figure 2, we normalize $A$ by their norms
$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}$ (10)
3.2 Linearity of Trajectories
Linearity in intermediate embeddings is quantified with the line-shape score (LSS), defined by Gai & Zhang (2021) as
$\text{LSS}_{i}^{0,\ldots,L}=\frac{L}{\left|\left|\tilde{x}^{L}_{i}-\tilde{x}^{0}_{i}\right|\right|_{2}},$ (11)
where $\tilde{x}_{i}^{0}=x_{i}^{0}$, i.e., the input embeddings passed to the LLM, and $\tilde{x}_{i}^{l}$ is defined recursively as
$\tilde{x}^{l}_{i}=\tilde{x}^{l-1}_{i}+\frac{x^{l}_{i}-x^{l-1}_{i}}{\left|\left|x^{l}_{i}-x^{l-1}_{i}\right|\right|_{2}}\ \ \mbox{for}\ \ l=1,\ldots,L.$ (12)
Note that $\text{LSS}\geq 1$, with $\text{LSS}=1$ if and only if the intermediate representations $x_{i}^{0},\ldots,x_{i}^{L}$ form a co-linear trajectory.
3.3 Layer-Wise Exponential Growth
We measure the presence of exponential spacing (expodistance) of the hidden trajectories. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can estimate,
$\|x_{i}^{l}\|$ $\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ (13)
for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$. We quantify the validity of the representation in Equation 13 by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
$\alpha_{i}^{l}\approx\ln\left(\frac{\|x_{i}^{l}\|}{\|x_{i}^{l-1}\|}\right),$ (14)
for each layer $l$ and token $i$. Under exponential growth, it is expected that $\alpha_{i}^{l}$ is independent of depth. We therefore denote the expodistance (ED) of the trajectory of the $i^{th}$ token of a given sequence by
$\text{ED}_{i}=\frac{\text{Var}_{l}\alpha_{i}^{l}}{(\text{Avg}_{l}\alpha_{i}^{l})^{2}}.$ (15)
This measurement is motivated by the parametrization discussed in Appendix A.4, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 17.
4 Evaluation
4.1 Suite of Large Language Models
Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations. These models, provided through HuggingFace (Wolf et al., 2020), vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens. Moreover, we analyze the dynamics of each measurement throughout training by deploying the Pythia Scaling Suite (Biderman et al., 2023). A summary of the models under consideration is presented in Table 1 of Appendix A.1 and further details in Appendix A.5.
4.2 Prompt Data
We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019). This data sets the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
5 Results
5.1 Coupling of Jacobians Across Depth
In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots. This is consistently observed across various LLMs considered. On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no coupling of Jacobians across different depths. There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors. This, in addition to Figure 2, suggests that coupling across depth emerges through training.
5.2 Coupling of Jacobians Across Tokens
We also analyze the coupling of singular vectors of the Jacobians across tokens. First, we observe coupling between Jacobians where the input and output tokens are the same, that is, between $J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$ (Figure LABEL:fig:coupling_inoutsame). This suggests that a token’s interaction on its own trajectory is coupled with that of other tokens. We also consider the context tokens’ interaction on a trajectory when fixing the input token $t_{1}$ (Figure LABEL:fig:coupling_fixedin), through $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$, and when fixing the output token $t_{2}$ (Figure LABEL:fig:coupling_fixedout), with $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$. Although generally present, the coupling in these two cases are less consistent, with variation in the strength of coupling between different token pairs. Again in untrained models (Figure LABEL:fig:coupling_inoutsame top row), there is a lack coupling of Jacobians across tokens.
5.3 Emergence of Coupling with Training
Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure LABEL:fig:paris_all_8to16 ) and across tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout). Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 12B, 6.9B, and 1B (Biderman et al., 2023) (Figure 2), and observe that coupling is generally low at initialization and increases significantly during early training steps.
5.4 Correlation with Generalization
For each LLM, we measure the average coupling across depth(which we define to be negative mis-coupling) across prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $m(A_{ll^{\prime},K}^{n})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$. We plot the coupling values against the benchmark scores across several LLMs (Figure 1). Our results reveal a positive correlation between coupling and performance benchmark scores, and is more significant than the relationship between other significant model hyperparameters (Figure LABEL:fig:correlation_hyperparams). This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
5.5 Regularity in Hidden Trajectories
We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs. The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure 13). Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training). Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
Figure 13: Linearity of Trajectories. The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 5.5), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. Models are sorted by number of parameters.
5.6 Layer-Wise Exponential Growth
Across all LLMs considered, the hidden trajectories exhibit exponential growth that emerges with training (Figure 14). The majority of the models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. The expodistance measurements show little dependence on model parameters, and take on consistently low values across all featured LLMs (Figure 14). In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 22). Under certain assumptions, exponential spacing is justified in Section 6.2.
Figure 14: Coefficient of variation of layer-wise exponent. The figure presents the coefficient of variation of the layer-wise exponent (Section 3.3) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for trained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale.
6 Discussion
6.1 Correlation with Generalization
As discussed in Section 5.4, stronger coupling across Jacobian singular vectors is related to improved generalization (Figure 1). This relationship may assist in evaluating the training of transformers. In particular, this finding suggests that the development of training methods to amplify coupling across transformer blocks may lead to favourable model performance.
6.2 Emergence of Exponential Spacing with Coupling
The emergence of exponential spacing in many LLMs can be viewed as a result of coupling under some assumptions on the homogeneity of the singular values across depth. Strong coupling suggests that the linearizations of hidden trajectories satisfy a linear difference equation (see Appendix A.4), given by
$x^{l+1}-x^{l}=Ax^{l}$
Considering the singular value decomposition of $A$ in Equation 18 as $A=U\Sigma V^{T}$, and assuming that there is coupling ($U=V$), the difference equation can be diagonalized in the same basis across depth. Further assuming that the top singular values of the Jacobian blocks display low variation across depth, (Figure LABEL:fig:svalues_trained), we expect
$||x^{l}||\approx e^{\sigma l}\,||x^{0}||,$ (16)
where $\sigma$ is the top singular value. Since we assume $\sigma$ is independent of depth, the above norm is expected to be exponential in $l$, and offers an explanation to the measurements in Figure 14.
7 Related Work
Residual Networks. ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021). The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment. Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023). We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023). Our measurements quantify the emergence with training of linearity and spacing, including at different depths, which was previously observed mostly qualitatively. Further comparison is included in Appendix A.3.
Neural ODE. Introduced by (Chen et al., 2018), Neural ODE view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE. (Zhong et al., 2022; Li et al., 2021) studied neural ODE specific perspective to transformers and other LLM architectures. In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
Information Processing in LLMs. Previously, Katz & Belinkov (2023); Bietti et al. (2023) have described memory and semantic flow in intermediate states of transformers while van Aken et al. (2019) have examined BERT hidden trajectories. Investigation of LLMs for reasoning tasks and interpretability remains an important research focus (Huang & Chang, 2023).
In-context learning. Previous works have observed that LLMs are able to perform in context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024): the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context. Their results consider a single layer of trained self-attention as implementing a single step of a gradient-descent-like learning rules across transformers depth, minimizing the MSE of a linear model
$x_{\text{final}}=\min_{x}||Ax-b||^{2}$
A gradient-descent learning rule would cause depth-wise updates to representations of the form
$x_{t+1}=x_{t}+\epsilon A^{T}(Ax_{t}-b)=(I+\epsilon A^{T}A)x_{t}-{\epsilon}A^{T}b$
We should expect similarity in the matrices $I+{\epsilon}A^{T}A$ across depth, which is suggested by coupling.
Hidden Representation Dynamics. Previous studies (Gai & Zhang, 2021; Haber & Ruthotto, 2017; Ee, 2017), have focused on interpreting deep neural networks from a dynamical systems perspective. Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space (Gai & Zhang, 2021). Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts (Nam et al., 2023). For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
Structure in Hidden Representations. Emergence of regularity in last-layer representations during training was noticed in Neural Collapse (Papyan et al., 2020). Subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023). Within LLMs, recent works have identified uniform structure in token means (Wu & Papyan, 2024), and low dimensional structure across hidden trajectories (Sarfati et al., 2024). Our work studies the representations of all layers in LLMs in a local sense through the interactions of tokens through their Jacobians.
8 Limitations
A limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them. Several of these models are trained independently of each other, without carefully controlling for different experimental parameters. Consequently, direct comparisons between models become challenging, and so determining the cause of varying levels of regularity between LLMs becomes difficult. Nonetheless, this makes the consistent emergence of these properties interesting for further study.
9 Conclusion
Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions. Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations. We additionally observe and quantify the emergence of linearity and exponential spacing with training in the embedding trajectories of LLMs. Notably, we establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization. These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
Appendix A Appendix and Supplementary material
A.1 Suite of Large Language Models and Prompt Data
We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
Base models.
Falcon (40B, 7B) (Almazrouei et al., 2023), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (6.9B).(Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
Fine-tuned models.
CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7.3B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
Datasets.
ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), WinoGrande (Sakaguchi et al., 2019).
Table 1: LLMs featured in the experiments throughout paper. Included in the table is the parameter budget of each model, the embedding dimension, the number of training tokens, and the Open LLM leaderboard (Beeching et al., 2023) benchmark score.
Model Param. Layers ($L$) Dim. ($d_{\text{model}}$) Tokens Score Pos.Encoding
Llama-3 70 B 80 8192 15 T RoPE
8 B 32 4096 15 T 62.35 RoPE
Llama-2 70 B 80 8192 2 T 66.05 RoPE
13 B 40 5120 2 T 55.69 RoPE
7 B 32 4096 2 T 50.97 RoPE
Codellama 34 B 48 8192 2 T 55.33 RoPE
13 B 40 5120 2 T 45.82 RoPE
7 B 32 4096 2 T 39.81 RoPE
Codellama (IT) 34 B 48 8192 2 T 43.0 RoPE
13 B 40 5120 2 T 37.52 RoPE
7 B 32 4096 2 T 40.05 RoPE
Orca 13 B 40 5120 2 T 58.64 RoPE
7 B 32 4096 2 T 54.55 RoPE
Falcon 40 B 60 8192 1 T 58.07 RoPE
7 B 32 4544 1.5 T 44.17 RoPE
Falcon (IT) 40 B 60 8192 1 T 43.26 RoPE
7 B 32 4544 1.5 T RoPE
MPT 30 B 48 7168 1 T 66.98 ALiBi
7 B 32 4096 1 T 56.83 ALiBi
Phi 2 B 32 2560 1.4 T 61.33 RoPE
1.5 B 24 2048 150 B 47.69 RoPE
1 B 24 2048 54 B RoPE
Mistral-v0.1 7.3 B 32 4096 60.97 ALiBi
Mistral-v0.1 (IT) 7.3 B 32 4096 54.96 ALiBi
Mistral-v0.2 (IT) 7.3 B 32 4096 65.71 ALiBi
Gemma 7 B 28 3072 6 T 64.29 RoPE
2 B 18 2048 6 T 42.75 RoPE
Gemma-1.1 7 B 28 3072 6 T 30.0 RoPE
2 B 18 2048 6 T 60.09 RoPE
Codegemma 7 B 28 3072 6 T 56.73 RoPE
2 B 18 2048 6 T 32.19 RoPE
Neo 20 B 44 6144 41.69 RoPE
2.7 B 32 2560 0.42 36.20 Sine
Pythia 12 B 36 5120 0.3 T 58.9 RoPE
6.9 B 32 4096 0.3 T 39.30 RoPE
2.8 B 32 2560 0.3 T RoPE
1.4 B 24 2048 0.3 T RoPE
1 B 16 2048 0.3 T RoPE
GPT-2 1.5 B 48 1600 34.12 Sine
774 M 36 1280 32.07 Sine
355 M 24 1024 29.87 Sine
117 M 12 768 28.53 Sine
A.2 Additional Metrics
A.2.1 Visualization of Trajectories with PCA
Each token, with initial embedding $x_{i}^{0}$, forms a trajectory $x_{i}^{0},x_{i}^{1},\ldots,x_{i}^{L}$ as it passes through the $L$ transformer blocks. The dynamics in high-dimensional space are visualized through a 2-dimensional principal component (PC) projection, $\text{PC}_{L}$, fitted to the last layer embeddings $X^{L}=(x_{1}^{L},x_{2}^{L},\ldots x_{n}^{L})$. The projected embeddings, $\text{PC}_{L}(x_{i}^{0}),\,\text{PC}_{L}(x_{i}^{1}),\,\ldots,\,\text{PC}_{L}(x_{i}^{L})$, are plotted for each of the $i=1,\ldots,n$ trajectories.
A.3 Comparison to ResNets
Coupling. Our results complement and build upon those of (Li & Papyan, 2023), who have observed the coupling of singular vectors of Residual Jacobians in classification ResNets. We observe coupling across depth in a wide range of LLMs, where Jacobians are evaluated at the sequence of embeddings, with respect to the current token, whereas in RA Jacobians are evaluated at a single representation. Additionally, with transformers we may analyze coupling not only across depth but also across tokens. We observe coupling in LLMs across tokens in a variety of ways. Further, we consider and analyze the relationship between coupling and generalization.
Linearity and Equidistance. Linearity in hidden trajectories, as observed in ResNets Li & Papyan (2023); Gai & Zhang (2021), also emerges with training in LLMs. The mean LSS value among the evaluated LLMs is $4.24$ (Figure 13), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension). In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories. In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 14). We quantify this spacing through a low coefficient of variation, displaying the presence of exponential growth in token trajectories. In both classifier ResNets and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
Rank of Jacobians. Li & Papyan (2023) show that residual Jacobians have rank at most $C$, the number of classes. This analogous result automatically holds for LLMs since the vocabulary size is significantly greater than the embedding dimension of the transformer blocks.
Singular Value Scaling. Li & Papyan (2023) observe that top singular values of residual Jacobians scale inversely with depth. In trained LLMs, however, top singular values do not show a consistent depth scaling across models (Figure LABEL:fig:svalues_trained), notably differing from classification ResNets. In addition, the distribution of singular values at each layer varies significantly between models. Singular value scaling is more present in untrained transformers (FigureLABEL:fig:svalues_untrained), likely caused by additional layer normalizations in residual blocks (Section 4.1).
A.4 Dynamical Motivation
The equality of top left and right singular vectors suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
$x^{\prime}=Ax$ (17)
Its solution is given by
$x(t)=\exp(tA)x(0)=\lim_{L\to\infty}\left(I+\frac{At}{L}\right)^{L}x(0)$ (18)
where the product is repeated $L$ times. Expanding the brackets shows that $x(t)$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. (2016) which views ResNets as
$x(t)=\left(I+\frac{tA_{1}}{L}\right)\left(I+\frac{tA_{2}}{L}\right)\dots\left(I+\frac{tA_{L}}{L}\right)x(0)$ (19)
However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{i}$ matrices. The coupling phenomenon suggests ResNets as implementing the simpler system
$x(t)=\left(I+\frac{tA}{L}\right)\left(I+\frac{tA}{L}\right)\dots\left(I+\frac{tA}{L}\right)x(0),$ (20)
where all the $A_{i}$ matrices are aligned. One benefit of this interpretation is that, as $L\to\infty$, we can write $x(t)$ in a simple closed form, namely $x(t)=\exp(tA)x(0)$.
Borrowing from the dynamical perspective, we quantify the similarity of hidden trajectories to the evolution of a linear ODE, in order to detect the emergence of a simple linearization to representations.
A.5 Experimental and Implementation Details
The source code used to produce the results reported in this experiment has been included as supplemental material. Models with varying parameter sizes are loaded on GPUs with appropriate memory requirements: NVIDIA A40 ($n_{\text{param}}\geq 40B$), NVIDIA Quadro RTX 6000 for Gemma variants and when ($40B>n_{\text{param}}>13B$), and NVIDIA Tesla T4 when ($13B\geq n_{\text{param}}$) except Gemma variants. 1,200 prompts from the OpenLLM leaderboard were evaluated in variable batch sizes were queued on a SLURM cluster, with appropriate adjustments depending on the memory required to load the LLM.
  • $13B\geq n_{\text{param}}$: 100 prompts per batch, except Gemma variants, which used 25 prompts per batch. The larger memory requirement for Gemma variants is likely due to the much larger vocabulary size in the model.
  • $40B>n_{\text{param}}>13B$: 10 prompts per batch, except NeoX 20B which used 100 prompts per batch.
  • $n_{\text{param}}\geq 40B$: 50 prompts per batch.
Due to the high memory requirement for computing Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized. Additionally, due to compute restrictions, alignment of singular vectors of Jacobians was computed on a smaller subset of the 1,200 prompts.
A.6 Additional Experimental Results
Figure 15: Hidden Trajectories in LLMs. Principal components of the trajectories of the hidden representations through various LLMs (columns, decreasing in model size, see Table 1) in the prompt: What is the capital of France? The capital is. Top row: all layers. Middle Row: layers in shallower transformer blocks (layers specified above plot). Bottom Row: layers in deeper transformer blocks (layers specified above plot). Trajectories of each input token (last token ‘is’ is plotted in black) are plotted in latent space, visualized with a 2-dimension principal component projection. Representations proceed in distinct outward directions, especially in the second half of transformer blocks (lower row) during which the norm of representations increases, with possible abrupt change in the last layer (outer points in upper row). A clear direction of movement is visible in each token trajectory.
Figure 21: Coefficient of variation of layer-wise equidistance. Variation of layer-wise equidistance (Section 3.3) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 4.2) on a suite of untrained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. The models are sorted by increasing benchmark performance.
Figure 22: Coefficient of variation of layer-wise expodistance. Variation of layer-wise expodistance (Section 3.3) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 4.2) on a suite of untrained LLMs (Appendix A.1).
Figure 23: Various Measurements of Representations. All measurements were made on 100 prompts taken from the WikiText 2 datsets. (Row 1) Norm of hidden representations as a function of layer depth. (Row 2) Line Shape Score (LSS) of the hidden trajectories as a function of layer depth. (Row 3) Mean equidistance of contiguous hidden trajectories as a function of depth. (Row 4) Entropy of logit vectors as a function of depth. Noted in most plots is a line where the behaviour of the measurement drastically changes.
Figure 24: Linearity Emerges with Training. Two plots displaying the evolution of the linearity of the token trajectories through training. (Left) The LSS as a function of training checkpoint for the variants of the Pythia Scaling Suite Biderman et al. (2023). Here, the LSS is measured over each entire prompt. (Right) The mean LSS as a function of layer depth measured at various checkpoints throughout the Pythia 12B model. Here, the LSS is computed on a window of layers of width $11$, centred at the value given by the x-axis.
Figure 32: LSS sorted by LLM parameters.